OpenStudy (lazyboy):

Given the function, Y = 3X, what are the Y values when X=0, X=1/3, and X = 10? Question 1 answers 0, 9, 30 3, 10/3, 13 0 , 9, 10 0, 1, 30

terenzreignz (terenzreignz):

Finally, no more sequences :D Just replace x with the given values, 0, 1/3, and 10

OpenStudy (anonymous):

did you replace \(x\) by the numbers to see what you get?

OpenStudy (lazyboy):

I'm on it

OpenStudy (anonymous):

i.e compute \[3\times 0\] \[3\times \frac{1}{3}\] \[3\times 10\]

OpenStudy (lazyboy):

still don't get it :(

terenzreignz (terenzreignz):

satellite's answer was direct. Just evaluate those products :)

OpenStudy (lazyboy):

So D

terenzreignz (terenzreignz):

Yeap. :)